You want to have $2.5 million in real dollars in an account when you retire in 40 years. The nominal return on your investment is 10.3 percent and the inflation rate is 3.7 percent.
 
What real amount must you deposit each year to achieve your goal? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Real rate=(1+nominal rate)/(1+inflation rate)-1

=(1+0.103)/(1+0.037)-1

=(1.103/1.037)-1

0.06364513(Approx)

Future value of annuity=Annuity[(1+rate)^time period-1]/rate

2,500,000=Annuity[(1+0.06364513)^40-1]/0.06364513

2,500,000=Annuity*169.6858857

Annuity=2,500,000/169.6858857

which is equal to

=$14733.11(Approx).